# Absolute abstraction: a renormalisation group approach

### v4

@@ Abstract
Abstraction refers to the ability to represent general concepts from sensory input or from raw data. It is well known that representations which are increasingly independent of details arise when raw data is processed in a hierarchy of deeper and deeper layers – both in-silico [1] and in-vivo [2]. Here we argue that universal representations emerge when the process of removing irrelevant details is iterated indefinitely on a universe of data that simultaneously expands in variety, that is when increasing depth is combined with expanding breadth [^1] We distinguish breadth from width, a term commonly used in the literature to denote the number of variables in different layers.. These universal representations encode the concept of absolute abstraction.
We address this in a minimal setting of unsupervised learning from static data (e.g. images). We shall focus on the internal representation [^2] A representation in this paper is a probability distribution over a set of binary variables. of the data in a deep layer of a learning machine and on how it adapts when the training data expands in scope. This process can be cast within a renormalisation group (RG) framework that allows us to identify the fixed point of the RG transformation with the ultimate outcome of this process when it is repeated indefinitely. The analogy between the process of learning broader and broader data in deeper and deeper leyers and the RG in statistical physics [3] (see Fig. 1 A) is based on the observation that higher order features in learning are akin to large scale properties in statistical models, which are those that the RG singles out. This idea is corroborated by similarities between coarse graining and the evolution of representations with depth [4, 5].
The fact that representations converge to a universal, data-independent, distribution is consistent with the expectation that a representation encompassing a broader universe of circumstances should be more abstract – i.e. data independent – than one describing only a limited domain. In this picture, all data-specific information is transferred to the parameters through which the activity of deep layers propagates to the visible layer. In other words, the fixed point of the RG describes how data is organised in the ideal limit of infinite depth and breadth, irrespective of what the data is about.
@@ 1 Literature review
The analogy between the processing of data in deeper and deeper layers and the renormalisation group (RG) in statistical physics [3] has been suggested [4, 5, 35] long ago. This analogy is suggestive because the RG is the theoretical tool to study critical phenomena [3] and both artificial and biological learning exhibit critical features [36, 37, 38, 39, 40, 41]. Koch et al. [5] have shown that successive training in deeper layers of neural networks performs an operation similar to coarse graining in the RG. Yet the RG transformation does not only involve coarse graining. It also involves rescaling, which is the operation by which the size of the coarse grained system is restored (see Fig. 1 A). We argue that this ingredient corresponds to a change in breadth, i.e. in the diversity in the input data or stimuli. It is indeed common sense that a representation encompassing a broader universe of circumstances should be more abstract than one describing only a limited domain. Likewise, we argue that representations in higher levels of the cognitive hierarchy, that integrate a broader set of stimuli, should be more abstract, i.e. independent of the data.
The emphasis on a data-independent characterisation of abstraction based only of statistical properties contrasts with the “tuning curve” approach [13, 42], in which levels of abstraction are assessed in terms of the features of the data – e.g. edges or faces – that a representation encodes. Likewise, although we implicitly assume that the representations we are interested in describe structured and learnable data, we shall refrain from assuming any specific structure like the ones proposed in Refs. [29, 43, 31].
@@ 2 The framework
In this paper, a representation is a probability distribution $p(\mathbf{s})$ over the configurations $\mathbf{s}$ of a hidden layer of a generative neural network, such as a Deep Belief Network (DBN) [44] or an auto-encoders [45]. Hence we focus on unsupervised learning and we assume that the network has been trained on some data $\hat{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ with a non-trivial structure. The representation $p(\mathbf{s})$ is the marginal distribution of activation of variables in the hidden layer induced by the distribution $p(\mathbf{x})$ of the data, i.e.
@@
We shall confine our discussion to the case where $\mathbf{s}_{1:n}=(s_{1},\ldots,s_{n})$ is a string of binary variables $s_{i}=0,1$. Otherwise, we shall keep our discussion as general as possible for the time being. More details on network architectures and the data will be given in Section 4 and Appendix C.2, where we shall test our predictions on specific models.
We may think of $s_{i}$ as indicator variables of abstract features: i.e. $s_{i}=1$ generates points with feature $i$ and $s_{i}=0$ does not. We shall also use, when needed, the notation $\mathbf{s}_{1:n}=(s_{1},\ldots,s_{n})$ to keep track of indices and $\mathbf{0}$ (or $\mathbf{0}_{1:n}$) to indicate the featureless state, i.e. the one with $s_{i}=0$ for all $i=1,\ldots,n$, which, as we shall see, describes most common objects.
@@ 3 Renormalization transformations
The renormalisation group (RG) [3] translates the process of focusing on large scale properties of statistical models into a mathematical formalism. As shown in Fig 1 A), this process is typically composed of two parts: i) coarse graining, whereby small scale details are eliminated and ii) rescaling in order to restore the original (length) scale. The combined effect of these two steps is a transformation from a probability distribution $p(\mathbf{s})$ to a different one $p^{\prime}=\Re[p]$ whose fixed points $p^{*}=\Re[p^{*}]$ describe scale invariant states. Fixed points are endowed with universal properties which make them independent of small scale details. In statistical physics, $p^{*}$ depends solely on few fundamental characteristics, such as symmetries and conservation laws, space dimension and dimensionality of the relevant variables (order parameters) [3]. In the context of learning, such universal distributions are natural candidates for an abstract representation, and the RG offers the ideal conceptual framework to search for them.
@@ 3.1 Zooming out
The representation capacity is limited by the number $n$ of available features and by the number of bits available to describe a single datapoint, which is measured by the entropy $H[\mathbf{s}]=-\sum_{\mathbf{s}}p(\mathbf{s})\log_{2}p(\mathbf{s})$. Within these limits, we assume that representations maximise a notion of informational efficiency. This means that features are used as parsimoniously as possible – as in sparse coding [46] – and that new data require the introduction of new large scale features which are in no way related to old ones. This reflects the principle that learning should venture into the unknown with no prejudice.
@@
    In this step, the representation is expanded to describe a wider universe of objects. The distribution of the new feature is independent of $\mathbf{s}_{1:n}$. This captures a genuine discovery process characterised by a large scale organisation (described by $s_{0}$) which cannot be described in terms of combinations of already known features.
@@
is the index of the most detailed active feature in $\mathbf{s}$, and $m_{\mathbf{0}}=0$. Interestingly, this distribution is related to the Hierarchical Feature model (HFM), introduced recently in Ref. [6], which is defined as
@@
where $g\geq 0$ is a parameter.
@@ 3.3 Significance of the results
Level of detail as sufficient statistics
The HFM was introduced in [6] as an efficient scaffold for data based on two principles. First, the HFM assumes that features are organised in a hierarchical scale of detail and it requires that the occurrence of a feature $s_{k}=1$ at level $k$ does not provide any information on whether lower order features are present or not. This means that, conditional on $s_{k}=1$, all lower order features are as random as possible, i.e. $H[\mathbf{s}_{1:k-1}|s_{k}=1]=k-1$ in bits. This requirement implies that the distribution $p({\mathbf{s}})$ should be a function of $m_{\mathbf{s}}$, as defined in Eq. (7). Indeed $m_{\mathbf{s}}$ quantifies the level of detail of objects associated to state $\mathbf{s}$. This is equivalent to requiring that the level of detail $m_{\mathbf{s}}$ is the only sufficient statistics of the distribution, in the sense that knowledge of the probability $p(m_{\mathbf{s}}=k)$ is sufficient to specify the full distribution $p({\mathbf{s}})$. If one further requires that just the average level of detail $\bar{m}=\langle m_{\mathbf{s}}\rangle$ is sufficient to determine $p({\mathbf{s}})$, i.e. that $p({\mathbf{s}})$ is a maximum entropy distribution, then this implies that the dependence of $p(\mathbf{s})$ on $m_{\mathbf{s}}$ should take an exponential form, which singles out Eq. (8) as the only possible solution.
@@
The maximum entropy requirement also coincides with demanding that $p(\mathbf{s})$ satisfies the principle of maximal relevance. The relevance is defined [9] as the entropy $H[E]$ of the coding cost $E_{\mathbf{s}}=-\log p(\mathbf{s})$ and is discussed in further detail in the Appendix. For our discussion, let it suffice to say that the relevance is a model-free measure of “meaningful” information content in a representation or in a data-set. Indeed, the internal representation of learning machines trained on non-trivial real data approach closely distributions of maximal relevance [10, 11, 9]. The principle of maximal relevance predicts that the number of states $\mathbf{s}$ with coding cost equal to $E$ should increase exponentially in $E$, i.e. $W(E)=W_{0}e^{\nu E}$ [39, 9]. The HFM satisfies this property with $\nu=\log 2$ because the number of states with $m_{\mathbf{s}}=k$ equals $2^{k-1}$ ($k>0$). Linearity of $E_{\mathbf{s}}$ with $m_{\mathbf{s}}$ implies that the number of bits required to describe a state $\mathbf{s}$ increases linearly with its level of detail.
@@
In the limit $n\to\infty$ the HFM features a phase transition at $g_{c}=\log 2$ between a random phase where the entropy $H[\mathbf{s}]$ and the average number of features $\langle m_{\mathbf{s}}\rangle$ are of order $n$ for $g<g_{c}$, and a “low temperature” phase where $h_{n}(\mathbf{s})$ is dominated by a finite number of states, with $H[\mathbf{s}]$ and $\langle m_{\mathbf{s}}\rangle$ attaining finite limits as $n\to\infty$). The distribution $p(\mathbf{s})$ for $g=g_{c}$ reproduces Zipf’s law, a statistical regularity that characterises many efficient representations, from language [50], to assemblies of neurons [38] and the immune system [51]. We refer to [6] for further discussion of the properties of the HFM.
The HFM is particularly intriguing for the peculiarity of its free energy landscape. The fact that the number of states $W(E)$ at energy $E$ grows exponentially with $E$ implies that the entropy $S(E)=k_{B}\log W(E)$ is linear in $E$. This means that also the free energy F(E)=ETS(E)F(E)=E-TS(E) is linear in $E$ and that there is a temperature for which $F(E)=F_{0}$ is constant. At this temperature, transitions from states of different energy require no work, in principle. These transitions, where either more details are added (increasing the level of detail $E$) or removed (decreasing $E$), are the natural building blocks of thought processes.
@@
For $g\leq g_{c}$ the constant $\alpha=1-2e^{-g}$ in the distribution $p^{*}$ of Eq. (6) becomes negative. So the marginal HFM Eq. (9) describes the fixed point of the transformation where the universe of the data expands only for $g\geq g_{c}$. Indeed, marginalising over the high order ones yields a mixture between the HFM and the uniform (maximum entropy) distribution (see Appendix B.2)
@@ 4 Empirical evidence in Deep Neural Networks
In this Section we test the ideas discussed in previous Sections on two architectures, Deep Belief Networks (DBNs) and auto-encoders (AE), trained on different datasets which are variants of the MNIST dataset of handwritten digits. A full account of architectures, algorithms and datasets is given in Appendix C.2. We retain the essential details in what follows.
@@ 4.1 Comparing internal representations with the HFM
First note that the two values $s_{i}=0$ or $1$ of each variable can be associated in two different ways to the values of $s_{i}^{\prime}$. Hence a state $\mathbf{s}$ defined by $n$ hidden binary variables, admits $2^{n}$ possible states $\mathbf{s}^{\prime}$, each consistent with the “gauge” $\mathbf{\tau}$ that defines the transformation $s_{i}\to s_{i}^{\prime}=\tau_{i}s_{i}+(1-\tau_{i})(1-s_{i})$ with $\tau_{i}=0,1$. In order to fix this gauge we set $\mathbf{\tau}$ such that the most frequently sampled state in each representation corresponds to the featureless state $\mathbf{s}^{\prime}=\mathbf{0}$, as for the HFM. In addition, while in the neural networks we shall study the variables $s_{i}$ are a priori equivalent, in the HFM they are not, because they are hierarchically organised. There are $n!$ possible ways to order the variables $\mathbf{s}_{1:n}^{\prime}$, that correspond to a different HFM. For a given permutation $\pi=(\pi(1),\ldots,\pi(n))$ of the integers $1,\ldots,n$, and a given gauge $\mathbf{\tau}$, the combined effect of these two operations defines the transformation
@@ 4.2 Deep Belief Networks
The results of extensive numerical experiments on DBNs are presented in Fig. 3, which displays the Kullback-Leibler divergence per node [^12] In the DBN we studied the number of hidden nodes $n_{\ell}$ vary substantially with depth $\ell$. We expect $D_{KL}(\hat{p}_{\ell}|\!|h_{n_{\ell}})$ to be proportional to $n_{\ell}$, which is why we show the Kullback-Leibler divergence per node. $D_{KL}(\hat{p}_{\ell}|\!|h_{n_{\ell}})/n_{\ell}$ in colour code (see bar on the left) as a function of breadth and depth (for layers [^13] We could not reliably estimate the distribution $p(\mathbf{s})$ for shallower layers. $\ell=5$ to $9$). We find that $D_{KL}(\hat{p}_{\ell}|\!|h_{n_{\ell}})/n_{\ell}$ generally decreases when both depth and breadth increase, as suggested by our RG approach. For a fixed depth, the internal representation initially approaches the HFM as breadth increases, and subsequently diverges from it. This suggests that both depth and breadth are necessary for abstraction to emerge, as claimed in the theoretical analysis of the previous Section.
@@
Panels a) and b) compare the same DBN trained for $10^{5}$ (a) or $10^{4}$ (b) epochs. As in other numerical experiments [52], we found that convergence to stable representations require long training times. We trained DBNs for a number of epochs ranging up to $10^{6}$. While convergence to stable distributions requires long training times (longer than $10^{4}$ epochs, as shown in c), training DBNs for too long time (e.g. $10^{6}$ epochs) leads to collapse of representations to distributions sharply peaked on one or two states. Panels a) and c) compare the same DBNs trained (for $10^{5}$ epochs) on the same datasets learned in a different order: in a) Fashion MNIST is introduced before EMNIST while in c) we did the opposite. This shows that the order in which datasets are learned matter: adding datasets which are more similar to those already learned results in a smoother approach to the HFM. The comparison between panels c) and d) – where the MNIST dataset is disaggregated in 5 parts which are added sequentially – further corroborates this conclusion. In the DBNs in d) the addition of the Cifar-10 dataset to ME3F (see caption) led to the collapse of the representation (which is why it was not shown). This is a manifestation of the limits of the representation capacity of these neural networks [53]. Panel e) shows that the fitted values of the parameter $g$ of the HFM decreases with breadth. This is consistent with the fact that in order to represent a wider universe of data, the internal representation needs to expand (indeed the entropy $H[\mathbf{s}]$ is a decreasing function of $g$). Results obtained fitting the marginal HFM Eq. (9) are consistent with this picture.
It has to be remarked that the representation $p(\mathbf{s})$ in all cases is rather far from the HFM, with a Kullback-Leibler divergence of $0.15$ bits per node in the best cases. This is also due to intrinsic limitations of our approach that cast doubts on whether larger DBNs and broader datasets would allow us to approach the HFM much further. We refer to the fact that our analysis is based on the representation in terms of binary variables $\mathbf{s}$, or $\mathbf{s}^{\prime}$ in Eq. (19). A closer look at representations of trained DBNs shows that $p(\mathbf{s})$ is characterised by several peaks [^14] Peaks can be defined through a simple heuristic or by computing the number of TAP solutions [54, 55].. Yet a metric based on the hamming distance is likely not the right one because the natural dynamics in DBNs is not a single variable dynamics but rather the one induced by the Montecarlo Markov chain (MCMC) with which states are sampled and the network is trained. A single MCMC step corresponds to transitions $\mathbf{s}\to\mathbf{s}^{\prime}$ that generally involve the update of more than one binary variable. We can probe the free energy in this metric by computing a transition matrix between labels in the following way: starting from a data point $\mathbf{x}$ of the dataset corresponding to a label $\alpha$ (e.g. a digit in MNIST) we sample one state $\mathbf{s}_{\ell}$ of the $\ell^{\rm th}$ layer. Then we propagate $\mathbf{s}_{\ell}$ to the visible layer and find the closest data point $\mathbf{x}^{\prime}$ and the associated label $\alpha^{\prime}$. Fig. 4 a) shows the structure of the transition matrix $T_{\alpha\to\alpha^{\prime}}^{(\ell)}$ which we obtain sampling many moves $\alpha\to\alpha^{\prime}$ as described above, for layers $\ell=1,4,7$ and $10$, in the MNIST dataset. In shallow layers, most likely all digits $\alpha$ “jump” to a data point corresponding to the same label $\alpha^{\prime}=\alpha$. In deep layers, transitions to other digits are much more likely. This suggests that shallow layers are characterised by many free energy minima, that become shallower and shallower in deeper layers, until they disappear in the deepest layers.
Figure 4: a) Colour plot of the transition matrix between digits (see text) generated by the representation in layer $\ell=1$ (top left), 44 (top right), 77 (bottom left) and $10$ (bottom right), of a DBN trained on MNIST. b) Distance $D(f_{\ell}^{\mathcal{D}},f_{\ell^{\prime}}^{\mathcal{D}^{\prime}})$ between the features $f_{\ell}^{\mathcal{D}}$ generated by states in layer $\ell$ of a DBN trained on dataset $\mathcal{D}$ and those ($f_{\ell^{\prime}}^{\mathcal{D}^{\prime}}$) generated by states in layer $\ell^{\prime}$ of a DBN trained on dataset $\mathcal{D}^{\prime}$. Features $f_{\ell}^{\mathcal{D}}=p(\mathbf{x}|\mathbf{s}_{\ell}=\mathbf{e}_{i})$ are defined as the activity patterns in the visible layer when only variable $i$ of layer $\ell$ is active. In b) the dataset $\mathcal{D}$ coincides with the MNIST digits 0 and 11 (M1), while $\mathcal{D}^{\prime}$ runs on all broader datasets $M2,\ldots,ME3F$.
If minima were to be identified with the object manifolds discussed, e.g., in Refs. [16], one would expect them to become sharper with depth according to the results of Cohen et al. [28]. Yet our analysis hints at the opposite conclusion. This might be related to the fact that while our analysis concentrates on unsupervised learning, Cohen et al. [28] discuss supervised learning, where labels induce non-trivial biases in the representations of deep layers. The multi-peak structure of shallow layers is also consistent with a scenario in which shallow layers rely on generic features to describe the data, and therefore semantic information [^15] In the present context, we call semantic information the mapping between the raw data and the labels. is contained in the activity $\mathbf{s}$. This is corroborated by the fact that the parameters of shallow layers do not change much when the network is successively trained on broader datasets, as shown in Fig. 4 b) (see caption). On the contrary, the parameters of deeper layers change considerably when the universe of data expands. This is consistent with semantic information being progressively stored in the “synapses” $p(\mathbf{x}|\mathbf{s}_{\ell})$, leaving the distribution $p(\mathbf{s})$ of activity patterns $\mathbf{s}$ more and more data invariant, as suggested by the RG analysis.

### v5

# arXiv:2407.01656v5[cs.LG] 03 Mar 2026

@@ Abstract
Abstraction refers to the ability to form general concepts from sensory input or raw data. It is well known that representations which are increasingly independent of details arise when raw data is processed in a hierarchy of deeper and deeper layers – both in-silico [1] and in-vivo [2]. Here we argue that universal representations emerge when the process of removing irrelevant details is iterated indefinitely on a universe of data that simultaneously expands in variety, that is when increasing depth is combined with expanding breadth [^1] We distinguish breadth from width, a term commonly used in the literature to denote the number of variables in different layers.. These universal representations encode the concept of absolute abstraction.
We address this in a minimal setting of unsupervised learning from static data (e.g. images). We shall focus on the internal representation [^2] A representation in this paper is a probability distribution over a set of binary variables. of the data in a deep layer of a learning machine and on how it adapts when the training data expands in scope. This process can be cast within a renormalisation group (RG) framework that allows us to identify the fixed point of the RG transformation with the ultimate outcome of this process when it is repeated indefinitely. The analogy between the process of learning broader and broader data in deeper and deeper layers and the RG in statistical physics [3] (see Fig. 1 A) is based on the observation that higher order features in learning are akin to large scale properties in statistical models, which are those that the RG singles out. This idea is further corroborated by similarities between coarse graining and the evolution of representations with depth [4, 5].
The fact that representations converge to a universal, data-independent, distribution aligns with the Platonic Representation Hypothesis [6], which is receiving more and more empirical support. This states that “Neural networks, trained with different objectives on different data and modalities, are converging to a shared statistical model of reality in their representation spaces” [6]. This hypothesis is consistent with the expectation that a representation encompassing a broader universe of circumstances should be more abstract – i.e. data independent – than one describing only a limited domain. In neural networks, this picture suggests that all data-specific information is transferred to the parameters through which the activity of deep layers propagates to the visible layer. In other words, the fixed point of the RG describes how data is organised in the ideal limit of infinite depth and breadth, irrespective of what the data is about.
@@ 1 Literature review
The analogy between the processing of data in deeper and deeper layers and the renormalisation group (RG) in statistical physics [3] has been suggested long ago [4, 5, 36]. This analogy is suggestive because the RG is the theoretical tool to study critical phenomena [3] and both artificial and biological learning exhibit critical features [37, 38, 39, 40, 41, 42]. Koch et al. [5] have shown that successive training in deeper layers of neural networks performs an operation similar to coarse graining in the RG. Yet the RG transformation does not only involve coarse graining. It also involves rescaling, which is the operation by which the size of the coarse grained system is restored (see Fig. 1 A). We argue that this ingredient corresponds to a change in breadth, i.e. in the diversity in the input data or stimuli. It is indeed common sense that a representation encompassing a broader universe of circumstances should be more abstract than one describing only a limited domain. Likewise, we argue that representations in higher levels of the cognitive hierarchy, that integrate a broader set of stimuli, should be more abstract, i.e. independent of the data.
The emphasis on a data-independent characterisation of abstraction based only of statistical properties contrasts with the “tuning curve” approach [14, 43], in which levels of abstraction are assessed in terms of the features of the data – e.g. edges or faces – that a representation encodes. Likewise, although we implicitly assume that the representations we are interested in describe structured and learnable data, we shall refrain from assuming any specific data structure like the ones proposed in Refs. [30, 44, 32].
@@ 2 The framework
In this paper, a representation is a probability distribution $p(\mathbf{s})$ over the configurations $\mathbf{s}$ of a hidden layer of a generative neural network, such as a Deep Belief Network (DBN) [45] or an auto-encoders [46]. Hence we focus on unsupervised learning and we assume that the network has been trained on some data $\hat{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ of $N$ data points, with a non-trivial structure. The representation $p(\mathbf{s})$ is the marginal distribution of activation of variables in the hidden layer induced by the distribution $p(\mathbf{x})$ of the data, i.e.
@@
We shall confine our discussion to the case where $\mathbf{s}=(s_{1},\ldots,s_{n})$ is a string of binary variables $s_{i}=0,1$. Otherwise, we shall keep our discussion as general as possible for the time being. More details on network architectures and the data will be given in Section 4 and Appendix C.2, where we shall test our predictions on specific models.
We may think of $s_{i}$ as indicator variables of abstract features: i.e. $s_{i}=1$ generates points with feature $i$ and $s_{i}=0$ does not. In order to keep track of indices, when needed, we shall use the notation $\mathbf{s}_{1:n}=(s_{1},\ldots,s_{n})$ for the state of the hidden layer, and $\mathbf{0}$ (or $\mathbf{0}_{1:n}$) to indicate the featureless state, i.e. the one with $s_{i}=0$ for all $i=1,\ldots,n$.
@@ 3 Renormalization transformations
The renormalisation group (RG) [3] translates the process of focusing on large scale properties of statistical models into a mathematical formalism. As shown in Fig 1 A), this process is typically composed of two parts: i) coarse graining, whereby small scale details are eliminated and ii) rescaling in order to restore the original (length) scale. The combined effect of these two steps is a transformation from a probability distribution $p(\mathbf{s})$ over the state $\mathbf{s}$ of the system, to a different one $p^{\prime}=\Re[p]$ whose fixed points $p^{*}=\Re[p^{*}]$ describe scale invariant states. Fixed points are endowed with universal properties which make them independent of small scale details. In statistical physics, $p^{*}$ depends solely on few fundamental characteristics, such as symmetries and conservation laws, space dimension and dimensionality of the relevant variables (order parameters) [3]. In the context of learning, such universal distributions are natural candidates for an abstract representation, and the RG offers the ideal conceptual framework to search for them.
@@ 3.1 Zooming out
The representation capacity is limited by the number $n$ of available features and by the number of bits available to describe a single datapoint, which is measured by the entropy $H[\mathbf{s}]=-\sum_{\mathbf{s}}p(\mathbf{s})\log_{2}p(\mathbf{s})$. Within these limits, we assume that representations maximise a notion of informational efficiency. This means that features are used as parsimoniously as possible, and that new data require the introduction of new large scale features which are in no way related to old ones. This reflects the principle that learning should venture into the unknown with no prejudice.
@@
    In this step, the representation is expanded to describe a wider universe of objects. The new feature $s_{0}$ has a distribution of maximal entropy and it is independent of $\mathbf{s}_{1:n}$. This captures a genuine discovery process characterised by a large scale organisation (described by $s_{0}$) which is maximally informative and that cannot be described in terms of combinations of already known features.
@@
is the index of the most detailed active feature in $\mathbf{s}$, that we shall call the level of detail, and $m_{\mathbf{0}}=0$. Interestingly, this distribution is related to the Hierarchical Feature model (HFM), introduced recently in Ref. [7], which is defined as
@@
where $g\geq 0$ is a parameter (see Appendix A.2 for a discussion of the HFM).
@@ 3.3 Significance of the results
Level of detail as sufficient statistics
The HFM was introduced in [7] as an efficient scaffold for data based on two principles. First, the HFM assumes that features are organised in a hierarchical scale of detail and it requires that the occurrence of a feature $s_{k}=1$ at level $k$ does not provide any information on whether lower order features are present or not. This means that, conditional on $s_{k}=1$, all lower order features are as random as possible, i.e. $H[\mathbf{s}_{1:k-1}|s_{k}=1]=k-1$ in bits. This requirement implies that the distribution $p({\mathbf{s}})$ should be a function of $m_{\mathbf{s}}$, as defined in Eq. (7), i.e. of the level of detail of objects associated to state $\mathbf{s}$. This is equivalent to requiring that the level of detail $m_{\mathbf{s}}$ is the only sufficient statistics of the distribution, in the sense that knowledge of the probability $p(m_{\mathbf{s}}=k)$ is sufficient to specify the full distribution $p({\mathbf{s}})$. If one further requires that just the average level of detail $\bar{m}=\langle m_{\mathbf{s}}\rangle$ is sufficient to determine $p({\mathbf{s}})$, i.e. that $p({\mathbf{s}})$ is a maximum entropy distribution, then this implies that the dependence of $p(\mathbf{s})$ on $m_{\mathbf{s}}$ should take an exponential form, which singles out Eq. (8) as the only possible solution.
@@
The maximum entropy requirement also coincides with demanding that $p(\mathbf{s})$ satisfies the principle of maximal relevance. The relevance is defined [10] as the entropy $H[E]$ of the coding cost $E_{\mathbf{s}}=-\log p(\mathbf{s})$ and is discussed in further detail in Appendix A.1. For our discussion, let it suffice to say that the relevance is a model-free measure of “meaningful” information content in a representation or in a data-set. Indeed, the internal representation of learning machines trained on non-trivial real data approach closely distributions of maximal relevance [11, 12, 10]. The principle of maximal relevance predicts that the number of states $\mathbf{s}$ with coding cost equal to $E$ should increase exponentially in $E$, i.e. $W(E)=W_{0}e^{\nu E}$ [40, 10]. The HFM satisfies this property with $\nu=\log 2$ because the number of states with $m_{\mathbf{s}}=k$ equals $2^{k-1}$ ($k>0$). Linearity of $E_{\mathbf{s}}$ with $m_{\mathbf{s}}$ implies that the number of bits required to describe a state $\mathbf{s}$ increases linearly with its level of detail.
@@
In the limit $n\to\infty$ the HFM features a phase transition at $g_{c}=\log 2$ between a random phase where the entropy $H[\mathbf{s}]$ and the average number of features $\langle m_{\mathbf{s}}\rangle$ are of order $n$ for $g<g_{c}$, and a “low temperature” phase where $h_{n}(\mathbf{s})$ is dominated by a finite number of states, with $H[\mathbf{s}]$ and $\langle m_{\mathbf{s}}\rangle$ attaining finite limits as $n\to\infty$. The distribution $h_{n}(\mathbf{s})$ for $g=g_{c}$ reproduces Zipf’s law, a statistical regularity that characterises many efficient representations, from language [50], to assemblies of neurons [39] and the immune system [51]. We refer to [7] for further discussion of the properties of the HFM.
The HFM is particularly intriguing for the peculiarity of its free energy landscape. The fact that the number of states $W(E)$ at energy $E$ grows exponentially with $E$ implies that the entropy $S(E)=k_{B}\log W(E)$ is linear in $E$. This means that also the free energy F(E)=ETS(E)F(E)=E-TS(E) is linear in $E$ and that there is a temperature for which $F(E)=F_{0}$ is constant. At this temperature, transitions from states of different energy require no work, in principle. These transitions, where either more details are added (increasing the level of detail $E$) or removed (decreasing $E$), are the natural building blocks of thought processes like generalization or specialization.
@@
For $g\leq g_{c}$ the constant $\alpha=1-2e^{-g}$ in the distribution $p^{*}$ of Eq. (6) becomes negative. So the marginal HFM Eq. (9) describes the fixed point of the transformation where the universe of the data expands only for $g\geq g_{c}$. Indeed, marginalising over the high order features yields a mixture between the HFM and the uniform (maximum entropy) distribution (see Appendix B.2)
@@ 4 Empirical evidence in Deep Neural Networks
In this Section we test the ideas discussed in previous Sections on two architectures, Deep Belief Networks (DBNs) and auto-encoders (AE), trained on different datasets which are variants of the MNIST dataset of handwritten digits. A full account of architectures, algorithms and datasets is given in Appendix C. We retain the essential details in what follows.
@@ 4.1 Comparing internal representations with the HFM
First note that the two values $s_{i}=0$ or $1$ of each variable can be associated in two different ways to the values of $s_{i}^{\prime}$. Hence a state $\mathbf{s}$ defined by $n$ hidden binary variables, admits $2^{n}$ possible states $\mathbf{s}^{\prime}$, each consistent with the “gauge” $\mathbf{\tau}$ that defines the transformation $s_{i}\to s_{i}^{\prime}=\tau_{i}s_{i}+(1-\tau_{i})(1-s_{i})$ with $\tau_{i}=0,1$. In order to fix this gauge we set $\mathbf{\tau}$ such that the most frequently sampled state in each representation corresponds to the featureless state $\mathbf{s}^{\prime}=\mathbf{0}$, as for the HFM. In addition, while in the neural networks the variables $s_{i}$ are a priori equivalent, in the HFM they are not, because they are hierarchically organised. There are $n!$ possible ways to order the variables $\mathbf{s}_{1:n}^{\prime}$, that correspond to a different HFM. For a given permutation $\pi=(\pi(1),\ldots,\pi(n))$ of the integers $1,\ldots,n$, and a given gauge $\mathbf{\tau}$, the combined effect of these two operations defines the transformation
@@ 4.2 Deep Belief Networks
The results of extensive numerical experiments on DBNs are presented in Fig. 3, which displays the Kullback-Leibler divergence per node [^12] In the DBN we studied the number of hidden nodes $n_{\ell}$ vary substantially with depth $\ell$. We expect $D_{KL}(\hat{p}_{\ell}|\!|h_{n_{\ell}})$ to be proportional to $n_{\ell}$, which is why we show the Kullback-Leibler divergence per node. $D_{KL}(\hat{p}_{\ell}|\!|h_{n_{\ell}})/n_{\ell}$ in colour code (see bar on the left) as a function of breadth and depth (for layers [^13] We could not reliably estimate the distribution $p(\mathbf{s})$ for shallower layers. $\ell=5$ to $9$). We found that $D_{KL}(\hat{p}_{\ell}|\!|h_{n_{\ell}})/n_{\ell}$ generally decreases when both depth and breadth increase, as suggested by the RG approach. For a fixed depth, the internal representation initially approaches the HFM as breadth increases and subsequently diverges from it. This suggests that both depth and breadth are necessary for abstraction to emerge, as claimed in the theoretical analysis of the previous Section.
@@
Fig. 3 suggests that this conclusion is robust with respect to varying the training time, the order in which datasets are learned and how they are segmented. Specifically, panels a) and b) compare the same DBN trained for $10^{5}$ (a) or $10^{4}$ (b) epochs. We found evidence that convergence to stable distributions requires long training times, as observed by Decelle and co-workers [52]. Consistently, we found that, although the approach to the HFM is observed in both cases, the convergence is smoother for longer training times (a) than for shorter times (b) [^14] Increasing further the training time (to $10^{6}$ epochs) led to the collapse of representations to distributions sharply peaked on one or two states.. Panels a) and c) compare the same DBNs trained (for $10^{5}$ epochs) on the same datasets learned in a different order: in a) Fashion MNIST is introduced before EMNIST while in c) we did the opposite. This comparison shows that, although both cases qualitatively agree with our thesis, the order in which datasets are learned matters: adding datasets which are more similar to those already learned results in a smoother approach to the HFM. The comparison between panels c) and d) – where the MNIST dataset is disaggregated in 5 parts which are added sequentially – further corroborates this conclusion [^15] In the DBNs in d) the addition of the Cifar-10 dataset to ME3F (see caption) led to the collapse of the representation (which is why it was not shown). This is a manifestation of the limits of the representation capacity of these neural networks [53].. Panel e) shows that the fitted values of the parameter $g$ of the HFM decrease with breadth. This is consistent with the fact that in order to represent a wider universe of data, the internal representation needs to expand (indeed the entropy $H[\mathbf{s}]$ is a decreasing function of $g$). Results obtained fitting the marginal HFM (Eq. 9) lead to similar conclusions.